a. The Tau Lepton has a proper lifetime of 0.3 picoseconds. What speed must tau particles have in order to travel an average distance of 1 mm.

b. Repeat this calculation with a distance of 107m.
Answer #1

(a)

We know that,

t = to / sqrt (1 - v2 / c2)

Speed of tau particles,

v = x / t

v = x / [to / sqrt (1 - v2 / c2)]

= (x / to) * sqrt (1 - v2 / c2)

= (10-3 / 0.3*10-12) sqrt (1 - v2 / (3*108)2)

v = 2.99*108 m/s

(b)

x = 107 m

v = x / [to / sqrt (1 - v2 / c2)]

= (x / to) * sqrt (1 - v2 / c2)

= (107 / 0.3*10-12) sqrt (1 - v2 / (3*108)2)

v = 2.99*1018 m/s